# Navigating Data Corruption in Machine Learning: Balancing Quality, Quantity, and Imputation Strategies

### v1

@@ 5 Effectiveness of Enlarging Dataset
Experiment Design
The aim of this section is to evaluate the effectiveness of enlarging the dataset and to quantify how much additional data is needed to offset the effects of data corruption. For the NLP-SL experiment, we recorded the performance of the model trained on datasets of different sizes and varying data corruption ratios. The variable “subset ratio” represents the proportion of the GLUE dataset used for fine-tuning. The results are shown in Figure 12 (a). For the Signal-RL task, we conducted experiments with different numbers of training episodes and data-missing ratios, as shown in Figure 12 (b).
Observations
As shown in the figure, as the dataset size increases, model performance converges. However, we observe that data corruption leads to a decline in model performance that cannot be fully recovered by increasing the sample size. When p is in the range [0, 0.4], the performance decline is nearly linear with respect to $p$ (Figure 12 (a)). For larger $p$, the model’s performance drops sharply to near zero (Figure 4).
This behavior is characteristic of an exponential function: for $e^{x}=1+x+\dots$, the linear term dominates when $x$ is small. Hence, we conclude that the performance drop increases approximately exponentially as the data corruption ratio increases. In addition, data corruption also hampers learning efficiency. To achieve the same level of performance (if achievable at all for a corrupted model), the number of samples—and therefore the training time—required increases exponentially with the data corruption level. This is illustrated by the dashed benchmark line in Figure 12 (b). Quantitative curves showing the relationship between data quality and the required amount of clean data provide practical insights into data collection strategies. Code and detailed results of this study are shared at Github repository.
Figure 12: Effectiveness of enlarging training dataset. (a) NLP-SL (b) Signal-RL. data corruption leads to a decline in model performance, which cannot be fully recovered by increasing the sample size. To achieve the same level of performance (if achievable by the corrupted model), the number of samples, and hence the training time, required increases exponentially with the data corruption level.
@@ 6 Conclusions
In this paper, we explored the impact of data corruption, including missing and noisy data, on deep learning performance across two distinct domains: supervised learning with NLP tasks (NLP-SL) and deep reinforcement learning for traffic signal optimization (Signal-RL). Our experiments aimed to provide insights into the relationship between data quality and model performance, the trade-offs of imputation strategies, and the effectiveness of increasing data quantity as a remedy for data corruption.
@@
    Diminishing Returns of Data Quality Improvement: Both NLP-SL and Signal-RL experiments revealed that model performance follows a diminishing return curve as data corruption decreases. The relationship is well-modeled by the exponential function $S=a(1-e^{-b(1-p)})$ , highlighting that additional clean data yields less utility once most patterns are recovered. This universal trend emphasizes the importance of balancing data quality and preprocessing efforts.
@@
    Noise vs. Missing Data: Our results demonstrate that noisy data is significantly more detrimental than missing data, leading to faster performance degradation and increased training instability. This was particularly evident in the Signal-RL task, where inserting noise caused substantial fluctuations in both training and final policy stability.
@@
    Imputation Trade-Offs: Imputation methods can restore critical information for missing data but introduce a trade-off by potentially adding noise. The decision to impute depends on the imputation accuracy and the corruption ratio. The “imputation advantage heatmap” revealed two regions: a “red corner” where accurate imputation significantly improves performance and a “blue edge” where noisy imputation harms the model. Tasks were categorized as “noise-sensitive” or “noise-insensitive” based on their decision boundaries, with RL tasks exhibiting greater sensitivity due to the compounding effects of sequential decision-making.
@@
    Enlarging Dataset as a Remedy: Increasing the dataset size partially mitigates the effects of data corruption but cannot fully recover the lost performance, especially under high noise levels. The analysis showed that the number of samples required to achieve a certain performance level increases exponentially with the corruption ratio, confirming the exponential nature of the trade-off between data quality and quantity.
@@
    Empirical Rule for Critical Data: An empirical rule emerged from the experiments: approximately 30% of the data is critical for determining the model’s performance, while the remaining 70% can be lost without substantial impact. This observation provides practical guidance for prioritizing efforts in data collection and preprocessing.
@@
Our findings offer valuable insights for machine learning practitioners and researchers. In resource-constrained environments, focusing on improving data quality for a small, critical portion of the dataset may yield higher returns than attempting to clean all data. Additionally, understanding the trade-offs of imputation strategies can inform the choice of preprocessing methods based on task sensitivity and corruption characteristics. For RL tasks, where noise compounds over sequential decisions, more robust strategies to handle corrupted observations are critical.
@@
This study opens up several avenues for future research. First, the observed patterns and empirical rules should be validated across broader datasets and additional machine learning tasks, such as computer vision and time-series forecasting. Second, developing adaptive imputation strategies that dynamically balance missing data recovery and noise introduction could further enhance model robustness. Finally, theoretical work on the relationship between information entropy, marginal utility, and model learning dynamics under corrupted data could deepen our understanding of these phenomena.
By addressing these challenges, we hope to advance the field’s ability to build robust machine learning models that perform reliably even in the presence of real-world data corruption.

### v2

@@ 5 Effectiveness of Enlarging Dataset
5.1 Experiment Design
The aim of this section is to evaluate the effectiveness of enlarging the dataset and to quantify how much additional data is needed to offset the effects of data corruption. For the NLP-SL experiment, the model is tested on datasets of different sizes and varying data corruption ratios. The variable “subset ratio” represents the proportion of the GLUE dataset used for fine-tuning. The results are shown in Figure 11 (a). The Signal-RL task experiments with different numbers of training episodes and data-missing ratios, as shown in Figure 11 (b).
5.2 Observations
As shown in the figure, as the dataset size increases, model performance converges. However, the results show that data corruption leads to a decline in model performance that cannot be fully recovered by increasing the sample size. When corruption ratio $p$ is in the range [0, 0.4], the performance decline is nearly linear with respect to $p$ (Figure 11 (a)). For larger $p$, the model’s performance drops sharply to near zero (Figure 3).
This behavior is characteristic of an exponential function: for $e^{x}=1+x+\dots$, the linear term dominates when $x$ is small. Hence, it’s concluded that the performance drop increases approximately exponentially as the data corruption ratio increases. In addition, data corruption also hampers learning efficiency. To achieve the same level of performance (if achievable at all for a corrupted model), the number of samples—and therefore the training time—required increases exponentially with the data corruption level. This is illustrated by the dashed benchmark line in Figure 11 (b). Quantitative curves showing the relationship between data quality and the required amount of data provide practical insights into data collection strategies.
Figure 11: Effectiveness of enlarging training dataset. (a) NLP-SL. The $x$-axis represents subuset ratio, and the $y$-axis indicates model score. (b) Signal-RL. The $x$-axis represents number of training episodes, and the $y$-axis indicates model score. Data corruption leads to a decline in model performance, which cannot be fully recovered by increasing the sample size. To achieve the same level of performance (if achievable by the corrupted model), the number of samples, and hence the training time, required increases exponentially with the data corruption level.
@@ 6 Conclusions
This study explored the impact of data corruption, including missing and noisy data, on deep learning performance across two distinct domains: supervised learning with NLP tasks (NLP-SL) and deep reinforcement learning for traffic signal optimization (Signal-RL). Our experiments aimed to provide insights into the relationship between data quality and model performance, the trade-offs of imputation strategies, and the effectiveness of increasing data quantity as a remedy for data corruption.
@@
    Diminishing Returns in Data Quality Improvement: Both NLP-SL and Signal-RL experiments revealed that model performance follows a diminishing return curve as data corruption decreases. The relationship between the model score $S$ and data corruption level $p$ is well-modeled by the function:
    $S=a(1-e^{-\lambda(1-p)})$
    where parameter $a=\frac{S_{0}}{1-e^{-\lambda}}$; and $S_{0}$ is the model score when corruption ratio $p$ = 0. This universal trend emphasizes the importance of balancing data quality and preprocessing efforts.
@@
    Data Noise is More Detrimental than Missing Data: Our results demonstrate that noisy data is significantly more detrimental than missing data, leading to faster performance degradation and increased training instability. This was particularly evident in reinforcement learning task, where inserting noise caused substantial fluctuations in both training and policy stability.
@@
    Trade-offs in Data Imputation: Imputation methods can restore critical information for missing data but introduce a trade-off by potentially adding noise, resulting in a trade-off. The decision to impute depends on the imputation accuracy, the corruption ratio, and the nature of the task. The imputation advantage heatmap highlights two key regions:
@@
    • Noise-insensitive tasks: These tasks exhibit gradual performance degradation, with decision boundaries on the heatmap that can be effectively modeled using a sigmoid curve.
    • Noise-sensitive tasks: These tasks exhibit sharp performance drops, with decision boundaries closely approximated by an exponential curve. This behavior is typical in deep reinforcement learning tasks. When only “general imputation” is available—as opposed to “exact imputation”—the sensitivity to noise tends to be further amplified.
@@
    Limits of Enlarging Datasets: Increasing the dataset size partially mitigates the effects of data corruption but cannot fully recover the lost performance, especially under high noisy levels. Enlarging datasets does not entirely offset the detrimental effects of noisy data. The analysis showed that the number of samples required to achieve a certain performance level increases exponentially with the corruption ratio, confirming the exponential nature of the trade-off between data quality and quantity.
  6. 6.
    Impact of Data Corruption on Learning Efficiency: Data missing hampers learning efficiency. To achieve the same performance level (if at all possible), the number of required samples—and hence training time—increases exponentially with the data missing level.
  7. 7.
    Empirical Rule on Data Importance: For traffic signal control tasks, approximately 30% of the data is critical for determining model performance, while the remaining 70% can be lost with minimal impact on performance. This observation provides practical guidance for prioritizing efforts in data collection and preprocessing. Note that the exact number may not apply to other tasks. And for many task this kind of screening is not feasible.
@@
The insights from this study provide direct implications for designing robust machine learning systems, benefiting both practitioners and researchers. Imputation strategies should be tailored to task-specific sensitivity, leveraging the identified boundary conditions between beneficial and harmful imputation. For noise-sensitive tasks, such as reinforcement learning, conservative imputation approaches and stricter data quality controls are essential to maintain model stability and performance. Practitioners should prioritize accurate data collection and preprocessing for the critical subset identified as having the most significant impact on performance. Specifically, in scenarios with limited resources, efforts should focus on minimizing noise in key data elements rather than indiscriminately expanding the dataset.
@@
Although this study evaluates two distinct tasks, the generalizability of its findings to other domains remains uncertain. The current study employs straightforward corruption methods, whereas real-world data corruption can be more complex, involving correlated or structured patterns of corruption not accounted for in the current experiments. Additionally, the artificial method of controlling imputation accuracy by inserting random noise provides useful insights into theoretical trade-offs but may not fully capture the nuances of practical imputation techniques. Moreover, although standard imputation methods were investigated, more advanced and sophisticated approaches remain underexplored, leaving their potential effectiveness unclear.
@@
This study opens up several avenues for future research. First, the observed patterns and empirical rules should be validated across broader datasets and additional machine learning tasks, such as computer vision and time-series forecasting. While the core principles—exponential performance decay, imputation decision boundaries—are hypothesized to generalize, their manifestations will likely depend on domain-specific factors. For instance, spatial correlations in CV or temporal dependencies in time-series data may influence how corruption impacts model performance. Notably, CV tasks may exhibit greater noise resilience compared to NLP, as pixel-level redundancies in images can mitigate localized corruption, whereas natural language is heavily compressed. Second, developing adaptive imputation strategies that dynamically balance missing data recovery and noise introduction could further enhance model robustness. Furthermore, numerous techniques have been developed for error detection, removal, and correction [5], investigating the effectiveness of these methods presents an interesting direction for future study. Finally, theoretical work on the relationship between information entropy, marginal utility, and model learning dynamics under corrupted data could deepen our understanding of these phenomena. By addressing these challenges, the hope is to advance the field’s ability to build robust machine learning models that perform reliably even in the presence of real-world data corruption.
Acknowledgment
This work was partly supported by research grant from Shanghai Baiyulan Talent Program Pujiang Project under grant number 24PJD115, Shanghai Yangpu District Postdoctoral Innovation & Practice Base Project.
Appendix A
This appendix contains the key details of the datasets, hyperparameter settings, and code implementation snippets referenced in the main text.
Table A1: Overview of model, dataset, and training configurations for NLP-SL and Signal-RL tasks
Task NLP-SL Signal-RL
Model type BERT DQN
Architecture & model description bert-base-uncased + classification_head
Vocab Size: 30,522 (WordPiece)
num_layers: 12
num_heads: 12
hidden_att: 768
hidden_ffn: 3072
hidden_dims: [256, 48]
State: road cell occupancy
Action: next phase for next 6 seconds
Action dim: 4
Step reward: $r_{t}=-\frac{q_{t}-80}{80}$
Stop speed threshold: 0.3 m/s
Datasets Pretraining: Wikitext, Bookcorpus
Finetuning: GLUE
50 simulation episodes
(1 Episode = 1h = 3600 steps)
Training config (Pretraining)
batch_size: 256
max_seq_len: 512
LR scheduler: linear warmup and decay
[…]
weight_decay: 0.01
(Seq: CoLA, SST2, MRPC, QQP, MNLI, QNLI, RTE, WNLI)
num_epochs: [5, 3, 5, 16, 3, 3, 3, 6, 2]
batch_size: [32, 64, 16, 16, 256, 256, 128, 8, 4]
lr: [3e-5, 3.5e-5, 3e-5, 3e-5, 5e-5, 5e-5, 5e-5, 2e-5, 1e-5]
weight_decay: 0.01
num_episodes: 50
batch_size: 256
LR scheduler: linear decay with plateau
[…]
Epsilon scheduler: linear decay with plateau
[…]
Discounting $\gamma$: 0.98
target_update: 10
buffer_size: 10000
Listing A1: Code for DQN model of Signal-RL
1class Qnet(nn.Module):
2 ’’’FFN with layer normalization, no dropout’’’
3 def __init__(self, state_dim, hidden_dims, action_dim, dropout_rate=0.1):
4 super().__init__()
5 self.fc1 = torch.nn.Linear(state_dim, hidden_dims[0])
6 self.ln1 = nn.LayerNorm(hidden_dims[0]) # LayerNorm after first hidden layer
7 self.fc2 = torch.nn.Linear(hidden_dims[0], hidden_dims[1])
8 self.fc3 = torch.nn.Linear(hidden_dims[1], action_dim)
9
10 def forward(self, x):
11 x = F.relu(self.ln1(self.fc1(x))) # Dimension: hidden_dims[0]
12 x = F.relu(self.fc2(x)) # Dimension: hidden_dims[2]
13 x = self.fc3(x) # Dimension: action_dim
14 return x
15
16class DQN:
17 """DQN class"""
18 def update(self, transition_dict):
19 # transition_dict is a dict of minibatch data
20 states = torch.tensor(transition_dict[’states’], dtype=torch.float).to(self.device)
21 actions = torch.tensor(transition_dict[’actions’], dtype=torch.int64).view(-1, 1).to(self.device)
22 rewards = torch.tensor(transition_dict[’rewards’], dtype=torch.float).view(-1, 1).to(self.device)
23 next_states = torch.tensor(transition_dict[’next_states’], dtype=torch.float).to(self.device)
24 dones = torch.tensor(transition_dict[’dones’], dtype=torch.float).view(-1, 1).to(self.device)
25 # compute target q
26 q_values = self.q_net(states).gather(1, actions)
27 max_action = self.q_net(next_states).max(1)[1].view(-1, 1) # Note: double DQN, use q_net to obtain next action
28 max_next_q_values = self.target_q_net(next_states).gather(1, max_action) # use target_q_net to obtain q values
29 q_targets = rewards + self.gamma * max_next_q_values
30 # compute loss
31 loss = F.smooth_l1_loss(q_values, q_targets) # Huber loss for more robust training
32 self.optimizer.zero_grad()
33 loss.backward()
34 self.optimizer.step()
35 # target network update
36 if self.count % self.target_update == 0:
37 self.target_q_net.load_state_dict(self.q_net.state_dict())
38 self.count += 1
Listing A2: Code for data corruption and artificial impuation (NLP-SL)
1def perturb_sentence(sentence):
2 words = sentence.split() # clean data
3 for i in range(len(words)):
4 if random.random() < p: # p: data missing level
5 original_word = words[i]
6 words[i] = tokenizer.unk_token # Replace with [UNK]
7 if imputation_method == ’insert_noise’: # artificial ’insert_noise’ imputation
8 if random.random() < q: # q is imputation noise level
9 words[i] = random.choice(whole_words) # Replace with a random whole word
10 else:
11 words[i] = original_word # Keep the original word
12 return " ".join(words)